OpenStudy (anonymous):

wrtie an equation of the line containing the given point and perpendicular to the given line. (0,8); 5x+2y=9 The equation of the line is y=?

OpenStudy (amistre64):

perps switch coeffitiients and negate one: 5x + 2y = n 2x -5y = n 2(0) -5(8) = n -40 = n 2x -5y = -40 y = (2/5) x +8